Calvin and Carolyn Coleman purchased a home in San Francisco for $375,000 on
October 1, 2017. Calvin obtained a job in Portland, Oregon, and on December 1, 2018,
the Colemans sold their home in San Francisco for $650,000. How much gain must the
Colemans recognize? (Use this to answer question below)

Assume that the Colemans in the preceding problem instead sold their home on December 1, 2018, for $800,000. How much gain must the Colemans recognize?
$425,000
$291,667
$0
$133,333
None of the above.
Answer #1

Home is used for qualified use for 14 months.

Maximum gain exclusion available = 500,000 X 14/24 = 291,667.

Realized gain = sale proceeds - basis = 800,000 - 375,000 = 425,000

Recognized gain = realized gain - gain exclusion = 425,000 - 291,667 = 133,333.

Correct option is $133,333.